Check my work Check My Work button is now enabledItem 2Item 2 8.33 points Item Skipped Speedy Auto Repairs uses a job-order costing system. The company’s direct materials consist of replacement parts installed in customer vehicles, and its direct labor consists of the mechanics’ hourly wages. Speedy’s overhead costs include various items, such as the shop manager’s salary, depreciation of equipment, utilities, insurance, and magazine subscriptions and refreshments for the waiting room. The company applies all of its overhead costs to jobs based on direct labor-hours. At the beginning of the year, it made the following estimates: Direct labor-hours required to support estimated output 40,000 Fixed overhead cost $ 640,000 Variable overhead cost per direct labor-hour $ 1.00 Required: 1. Compute the predetermined overhead rate. 2. During the year, Mr. Wilkes brought in his vehicle to replace his brakes, spark plugs, and tires. The following information was available with respect to his job: Direct materials $ 707 Direct labor cost $ 230 Direct labor-hours used 9 Compute Mr. Wilkes’ total job cost. 3. If Speedy establishes its selling prices using a markup percentage of 50% of its total job cost, then how much would it have charged Mr. Wilkes?
Answer #1

1) Predetermined Overhead rate = Estimated Fixed Overhead Cost/Estimated Direct Labor Hours

= 640,000/40,000 hrs = $16 per hour

2) Calculation of Mr. Wilkes Total Job Cost (Amounts in $)

Direct materials 707
Direct labor cost 230
Variable Overhead cost (9 hrs*$1) 9
Allocated Fixed Overhead Cost (9 hrs*$16) 144
Mr. Wilkes Total Job Cost 1,090

Therefore Mr. Wilkes total job cost is $1,090.

3) Markup charged from Mr. Wilkes = Total job cost*50%

= $1,090*50% = $545

Amount charged from Mr. Wilkes = Total Job Cost+Markup

= $1,090+$545 = $1,635

Therefore total amount charged from Mr. Wilkes would be $1,635.